OpenStudy (anonymous):

One pump can drain a pool in 10 minutes. When a second pump is also used, the pool only takes 4 minutes to drain. How long would it take the second pump to drain the pool if it were the only pump in use?

OpenStudy (anonymous):

yeah see..... 1/10 + b = 1/9 per minute. b = 1/90 per minute. 90 minutes to drain it complete... Hope it helps!!

OpenStudy (anonymous):

I was intended to give you an answer, but the similar question is explained in the site give below. If you don't understand ask me, I would be glad to provide assistance. check the below link for the steps for similar problem. http://www.algebra.com/algebra/homework/Rate-of-work-word-problems/Rate-of-work-word-problems.faq.question.230986.html